# What Determines the Maximum Mass of AGN-assisted Black Hole Mergers?

### v1

@@ 5 Conclusion
We carried out a set of AGN simulations to model AGN-assisted black hole mergers. In particular, we aimed to determined the maximum black hole mass expected to be observed from AGNs as a function of binary and AGN parameters (see Table 1). We examined, through multiple simulations, the dependence of the maximum mass on each of our model parameters, while keeping all other parameters at its fiducial value to be able to explore the 24-dimensional parameters space (see Fig. 12 and 10). Our conclusions are the following.
@@
    We find that a black hole mass of $\gtrsim 200$ M is only attainable for AGN lifetimes $\gtrsim 40$ Myr, independently of the other parameters. Therefore, the possible future detection of such a high-mass black hole in a merger, and the identification of the host AGN through electromagnetic followup, would give a unique measurement of the minimum lifetime of an AGN disk.
@@
    We similarly find the dependence of the AGN-assisted merger rate as a function of each of our model parameters (see Figs. 11 and 10), which can also be used to constrain the AGN scenario once it becomes sufficiently clear what subset of the observed black hole mergers are of AGN origin.

### v2

@@ 5 Conclusion
We carried out a set of AGN simulations to model AGN-assisted black hole mergers. In particular, we aimed to determined the maximum black hole mass expected to be observed from AGNs as a function of binary and AGN parameters (see Table 1). We examined, through multiple simulations, the dependence of the maximum mass on each of our model parameters, while keeping all other parameters at its fiducial value to be able to explore the 24-dimensional parameters space (see Fig. 7 and 5). Our conclusions are the following.
@@
    We found that a black hole mass of $\gtrsim 200$ M is only attainable for AGN lifetimes $\gtrsim 40$ Myr, independently of the other parameters. Therefore, the possible future detection of such a high-mass black hole in a merger, and the identification of the host AGN through electromagnetic followup, would give a unique measurement of the minimum lifetime of an AGN disk.
@@
    We similarly find the dependence of the AGN-assisted merger rate as a function of each of our model parameters (see Figs. 6 and 5), which can also be used to constrain the AGN scenario once it becomes sufficiently clear what subset of the observed black hole mergers are of AGN origin.